# Advances in Protein Representation Learning: Methods, Applications, and Future Directions

### v1

@@ Abstract
A protein’s molecular surface is a compact, smooth boundary formed by the outermost atoms, exhibiting distinct chemical and geometric patterns. As the primary interface for molecular interactions, the protein surface plays a crucial role in molecular recognition, ligand binding, and protein-protein interactions. Protein surface representation focuses on modeling this exterior, where these interactions predominantly occur, providing a functionally relevant perspective that complements residue-level and atomic-level representations.
@@
Several notable studies have advanced protein surface representations by leveraging 3D mesh-based approaches 43, 44, 45, 46, 47 and point cloud-based approaches 48, 49, 50, 51. In mesh-based models, the protein surface is represented as a triangulated mesh, where nodes correspond to discrete surface points, and edges define geometric relationships between neighboring points, preserving the overall topology of the surface. These models effectively capture geometric features such as curvature, solvent accessibility, and surface normals, as well as chemical properties like electrostatic potential and hydrophobicity. Due to their structured nature, mesh-based representations excel in tasks that require fine-grained surface characterization, such as functional site identification, enzyme active site modeling, and predicting molecular docking orientations. However, mesh-based approaches often rely on precomputed geometric and chemical features, making them computationally demanding and less adaptable for large-scale datasets.
@@
In contrast, point cloud-based approaches offer a more computationally efficient alternative by avoiding the need for precomputed features. These models operate directly on the raw set of atoms composing the protein, generating a point cloud representation of the surface on the fly. Unlike mesh-based methods, point cloud approaches learn task-specific geometric and chemical features dynamically and apply convolutional operators that approximate geodesic coordinates in the tangent space. This eliminates the need for an explicit surface mesh and significantly reduces memory usage and computational overhead, making point cloud-based representations well-suited for large-scale protein-ligand binding predictions, protein-protein interface modeling, and high-throughput interaction analysis.
4.4 4.4. Symmetry-Preservation and Equivariance in Structural Representation
@@
Incorporating three-dimensional (3D) structural information into PRL presents the challenge of ensuring models correctly handle spatial transformations such as rotations and translations. Models relying on raw 3D coordinates risk developing spurious dependencies on absolute positioning, leading to inconsistent predictions for identical structures in different orientations. This limitation is particularly problematic in tasks like protein folding, molecular docking, and binding site prediction, where relative spatial geometry governs molecular behavior. Symmetry-preserving and equivariant models addresses this issue by explicitly incorporating the inherent rotational and translational symmetries of protein structures, improving model performance, generalization, and computational efficiency. Such models ensure either invariance (i.e. output representation remains unchanged regardless of input transformations) or equivariance (i.e. output representation transforms accordingly to input transformations), preserving relative spatial relationships.
@@
Recent advancements in geometric deep learning have led to the development of rotation-equivariant 52, 53, 54 and permutation-equivariant graph neural networks 55, 56, which explicitly encode spatial symmetries, improving PRL 31, 32, 39, 35. By inherently preserving geometric properties, these models enhance the accuracy of predictions in tasks such as protein structure modeling and molecular interactions. Moreover, by eliminating the need for extensive data augmentation—such as generating rotated versions of protein structures—they reduce computational overhead while improving model robustness and efficiency.
@@
In parallel, invariant models address spatial symmetries by learning transformation-independent features, such as interatomic distances, bond angles, and backbone dihedral angles 50, 33, 46, 57. These geometric descriptors remain constant under rigid-body transformations, ensuring that identical protein structures in different orientations yield the same representation. For instance, atomic distances, bond angles, and dihedral angles capture essential spatial relationships while being inherently independent of absolute positioning. This approach simplifies learning, reduces computational complexity, and enhances generalization across diverse protein structures. By eliminating orientation and positional variance, invariant models streamline training while maintaining high predictive performance.
5 5. Multimodal-Based Approaches
@@
5.1 5.1. Integration of Sequence and Structure Representations
@@
    The first approach focuses on integrating sequence-derived information into structure-based models. In this method, PLMs are used to extract amino acid embeddings from sequences, which are then treated as node features in residue-level graphs 58, 59, 60, 61, 62. These graphs are constructed using structural data, capturing spatial dependencies between residues through edges defined by proximity, interaction types, or geometric constraints. This strategy leverages the contextual richness of PLMs while grounding the representations in the protein’s three-dimensional structure, resulting in more comprehensive and informative models.
  2. 2.
    The second approach takes the reverse direction by integrating structural information directly into the training of PLMs. Instead of relying solely on sequence data, this method embeds geometric and spatial properties into the PLM framework, allowing the model to process structural features alongside amino acid sequences 63, 64, 65, 66, 67. By incorporating structural cues, such as residue spatial proximity or geometric constraints, PLMs can generate richer, more biologically relevant representations that capture both the linear sequence and three-dimensional conformation of proteins.
  3. 3.
@@
    The third approach treats sequence and structural data as distinct modalities, encoding them separately using models tailored to each 68, 69, 57, 70, 71, 72, 42, 73. The resulting representations are then fused through techniques such as attention mechanisms, contrastive learning, or embedding concatenation. This multi-branch strategy allows models to learn complementary features from both modalities while preserving the unique contributions of each, leading to enhanced representation quality and generalizability.
5.2 5.2. Integration of Sequence and Functional Representations
@@
While Protein Language Models (PLMs) have advanced the understanding of sequence patterns, they often lack the explicit biological knowledge required to accurately capture protein function. Sequence data alone provides limited insight into a protein’s specific roles or its molecular interactions within cellular environments, which can hinder model performance in tasks such as protein function prediction, protein contact mapping, and protein-protein interaction analysis. Given that a protein’s structure fundamentally determines its function, models can be significantly improved by incorporating functional knowledge that highlights similarities among proteins with comparable shapes or sequences. Functional annotations provide comprehensive insights into a protein’s biological processes, molecular functions, and cellular components. Integrating such functional data into PLMs enriches protein representations, allowing models to more effectively generalize across diverse biological tasks and applications.
@@
To address the limitations of sequence-only models, functional data is typically treated as a separate modality. Researchers employ biomedical language models like BioBERT 74 and PubMedBERT 75, which are trained on large-scale biomedical literature and ontologies, to generate functional embeddings that capture rich contextual and domain-specific knowledge. These functional embeddings are then fused with sequence-derived representations using methods such as contrastive learning or joint embedding frameworks 76, 77, 78, 79. This multimodal integration enhances the representational capacity of PLMs, enabling more accurate and biologically informed predictions across a range of functional and interaction-based tasks.
5.3 5.3. Unified Representations of Sequence, Structure, and Function
@@
Unified representations aim to simultaneously integrate sequence, structural, and functional data into a cohesive modeling framework, providing a comprehensive view of protein characteristics. While sequence-based models capture evolutionary and contextual relationships, and structure-based models provide insights into three-dimensional conformation, integrating functional annotations adds critical biological context regarding a protein’s roles and interactions within cellular systems. By combining these modalities, unified representations offer a holistic understanding of proteins, enabling models to capture the complex interplay between a protein’s linear sequence, its folded structure, and its biological function.
@@
To achieve this, researchers typically adopt multi-branch architectures, where each modality—sequence, structure, and function—is encoded separately using specialized models. The outputs from these models are then fused using techniques such as attention mechanisms, contrastive learning, or cross-modal embedding fusion 80, 81. This approach ensures that the unique contributions of each modality are preserved while enabling the model to learn complementary features that enhance overall representation quality. By capturing the full complexity of protein biology, these multimodal approaches pave the way for more accurate, generalizable, and biologically meaningful models.
6 6. Complex-Based Approaches
@@
6.1 6.1. Protein-Ligand Complex Representation
@@
Protein-ligand complex representation aims to learn a unified representation that effectively captures the interactions between proteins and small molecules, enabling models to better understand binding mechanisms and affinity relationships. Rather than representing the entire protein structure, these approaches typically focus on the binding pocket, the localized region where molecular interactions occur. By concentrating on this functionally relevant site, models can more accurately capture the spatial, chemical, and sequence-dependent features that govern ligand binding and molecular recognition. Approaches for protein-ligand complex representation can be broadly classified into two following categories, differentiated by their use of 3D structural information.
@@
The first category follows a dual-encoder framework 82, 83, 84, 85, 86, 87, 88, where proteins and ligands are encoded separately, without explicitly modeling their spatial interactions. Protein representations are commonly derived from protein language models (PLMs) or graph-based encoders, while ligand representations are generated using SMILES-based transformers, molecular fingerprints, or graph neural networks (GNNs). Since these models do not require pre-existing 3D structural data for the full complex, they are particularly advantageous for high-throughput virtual screening and cases where structural information is unavailable. During fine-tuning, a lightweight interaction module is introduced to align and refine these representations for downstream tasks such as binding affinity prediction and molecular docking. This modular approach enhances scalability and facilitates transfer learning across diverse protein-ligand datasets.
@@
The second category focuses on fine-grained interaction modeling, where 3D structural information of the full protein-ligand complex is explicitly incorporated to learn atomic or residue-level interactions. Unlike dual-encoder approaches, which align representations post hoc, these methods directly encode spatial and chemical dependencies within the binding pocket, capturing fine-grained molecular interactions. The binding mechanism of protein-ligand complexes is highly intricate, involving a diverse array of non-covalent forces such as $\pi$-stacking, $\pi$-cation interactions, salt bridges, water bridges, hydrogen bonds, hydrophobic interactions, and halogen bonds 89. Previous studies 90, 91, 92 have demonstrated that explicitly modeling these atomic-level forces significantly improves predictive performance, underscoring the importance of 3D molecular force modeling in PRL.
6.2 6.2. Protein Complex Representation
@@
Protein complexes, consisting of two or more interacting polypeptide chains, are fundamental to biological processes such as signal transduction, immune response, enzymatic activity, and molecular assembly. These complexes exhibit significant diversity in size, stability, and interaction mechanisms, ranging from small, transient interactions to large, stable macromolecular structures such as ribosomes, chaperones, and antibody-antigen complexes. The representation of protein complexes closely follows the methodology used in protein-ligand complex modeling, where the goal is to learn a joint representation that captures both individual protein properties and interaction-specific features. Approaches to this problem can be broadly categorized based on whether they explicitly encode the full complex structure or rely on separate representations of individual proteins.
@@
Structure-free approaches, which encode individual proteins separately before aligning them in a shared representation space, are widely applied to various challenges in protein complex modeling 93, 94, 95. These approaches do not explicitly model the structural constraints of the full complex during encoding. Instead, interactions are inferred post hoc through mechanisms such as contrastive learning or learned scoring functions. Among the many tasks tackled by these approaches, protein complex structure prediction, determining the spatial arrangement of interacting polypeptides given only sequence or monomeric structural information, is one of the most critical 96, 97, 98.
@@
Structure-aware approaches, on the other hand, explicitly encode the entire 3D structure of the protein complex, incorporating spatial and physicochemical constraints to model binding interfaces and interaction dependencies 99, 100, 101. These methods leverage geometric deep learning, graph-based modeling, and energy-based optimization techniques to refine complex formation by capturing atomic-level interactions and physical constraints. Because these approaches explicitly encode the joint structural context of interacting proteins, they can provide more interpretable predictions, improve docking accuracy, and enhance the design of protein-protein interactions for therapeutic applications. However, they require high-quality structural data, which may not always be available, and must balance computational efficiency with accuracy when handling large macromolecular assemblies.
Table 1: Comparative Summary of Strengths and Limitations of PRL Approaches.
@@
7 7. Databases for Protein Representation Learning
Large-scale databases are the backbone of PRL, providing the essential sequence, structure, and functional data required for model training, validation, and benchmarking. This section introduces key databases widely used in PRL, categorized into sequence, structure, and protein function databases, as summarized in Table 2. These databases serve as fundamental resources for developing and evaluating protein representation models, enabling advancements in various biological and computational applications.

### v2

@@ 4.3. Protein Surface Representation
A protein’s molecular surface is a compact, smooth boundary formed by the outermost atoms, exhibiting distinct chemical and geometric patterns. As the primary interface for molecular interactions, the protein surface plays a crucial role in molecular recognition, ligand binding, and protein-protein interactions. The protein surface representation focuses on modeling this exterior, where these interactions predominantly occur, providing a functionally relevant perspective that complements residue-level and atomic-level representations.
@@
Several notable studies have advanced protein surface representations by using 3D mesh-based approaches (gainza2020deciphering, ; riahi2023surface, ; mallet2023atomsurf, ; randolph2024invariant, ; marchand2025targeting, ) and point cloud-based approaches (Sverrisson_2021_CVPR, ; liu2021octsurf, ; Srinivasan2022, ; sun2023dsr, ). In mesh-based models, the protein surface is represented as a triangulated mesh, where the nodes correspond to discrete surface points, and the edges define geometric relationships between neighboring points, preserving the overall topology of the surface. These models effectively capture geometric features such as curvature, solvent accessibility, and surface normals, as well as chemical properties such as electrostatic potential and hydrophobicity. Due to their structured nature, mesh-based representations excel in tasks that require fine-grained surface characterization, such as functional site identification, enzyme active site modeling, and molecular docking orientation prediction. However, mesh-based approaches often rely on precomputed geometric and chemical features, making them computationally demanding and less adaptable for large-scale datasets.
@@
In contrast, point-cloud-based approaches offer a more computationally efficient alternative by avoiding the need for precomputed features. These models operate directly on the raw set of atoms composing the protein, generating a point-cloud representation of the surface on the fly. Unlike mesh-based methods, point cloud approaches learn task-specific geometric and chemical features dynamically and apply convolutional operators that approximate geodesic coordinates in the tangent space. This eliminates the need for an explicit surface mesh and significantly reduces memory usage and computational overhead, making point cloud-based representations well-suited for large-scale protein-ligand binding predictions, protein-protein interface modeling, and high-throughput interaction analysis.
4.4. Symmetry-Preservation and Equivariance in Structural Representation
@@
Incorporating three-dimensional (3D) structural information into PRL presents the challenge of ensuring that models correctly handle spatial transformations such as rotations and translations. Models relying on raw 3D coordinates risk developing spurious dependencies on absolute positioning, leading to inconsistent predictions for identical structures in different orientations. This limitation is particularly problematic in tasks such as protein folding, molecular docking, and prediction of binding sites, where the relative spatial geometry governs the molecular behavior. Symmetry-preserving and equivariant models address this issue by explicitly incorporating the inherent rotational and translational symmetries of protein structures, improving model performance, generalization, and computational efficiency. Such models ensure either invariance (i.e. output representation remains unchanged regardless of input transformations) or equivariance (i.e. output representation transforms accordingly to input transformations), preserving relative spatial relationships.
@@
Recent advances in geometric deep learning have led to the development of rotation-equivariant (pmlr-v139-satorras21a, ; NEURIPS2020_15231a7c, ; NEURIPS2019_03573b32, ) and permutation-equivariant graph neural networks (maron2018invariant, ; 10.1063/1.5024797, ), which explicitly encode spatial symmetries, improving PRL (zhou2022lightweight, ; le2022representation, ; wu2023atomic, ; chen20233d, ). By inherently preserving geometric properties, these models enhance the accuracy of predictions in tasks such as protein structure modeling and molecular interactions. Moreover, by eliminating the need for extensive data augmentation, such as generating rotated versions of protein structures, they reduce computational overhead while improving model robustness and efficiency.
@@
In parallel, invariant models address spatial symmetries by learning transformation-independent features, such as interatomic distances, bond angles, and dihedral angles of the backbone (Srinivasan2022, ; guo2022self, ; randolph2024invariant, ; cheng2024zero, ). These geometric descriptors remain constant under rigid-body transformations, ensuring that identical protein structures in different orientations yield the same representation. For instance, atomic distances, bond angles, and dihedral angles capture essential spatial relationships while being inherently independent of absolute positioning. This approach simplifies learning, reduces computational complexity, and improves generalization across diverse protein structures. By eliminating orientation and positional variance, invariant models streamline training while maintaining high predictive performance.
5. Multimodal-Based Approaches
@@
5.1. Integration of Sequence and Structure Representations
@@
    The first approach focuses on the integration of sequence-derived information into structure-based models. In this method, PLMs are used to extract amino acid embeddings from sequences, which are then treated as node features on residue-level graphs (wang2022lm, ; gligorijevic2021structure, ; wu2023integration, ; blaabjerg2024ssemb, ; ahmed2025improved, ). These graphs are constructed using structural data, capturing spatial dependencies between residues through edges defined by proximity, interaction types, or geometric constraints. This strategy takes advantage of the contextual richness of PLMs while integrating the representations in the three-dimensional structure of the protein, resulting in more comprehensive and informative models.
  2. (2)
    The second approach takes the reverse direction by integrating structural information directly into the training of PLMs. Instead of relying solely on amino acid sequences, this method introduces structural data, such as residue distances, angles, or spatial neighbors, into the input or learning objective of the model (zheng2023structure, ; yang2023masked, ; su2024saprot, ; sun2024structure, ; li2025large, ). In some cases, models replace or augment traditional sequence tokens with residue-level structural tokens that describe the 3D context of each amino acid. These tokens, often derived from geometric relationships or learned structural alphabets, allow the model to process structural patterns using the same language-modeling techniques applied to sequences. By learning both the sequence and the structure at the token level, these models can generate more informative and biologically grounded representations.
  3. (3)
@@
    The third approach treats sequence and structural data as distinct modalities, encoding them separately using models tailored to each (lee2023pre, ; wang2023s, ; cheng2024zero, ; zhang2024pre, ; barton2024enhancing, ; ngo2024multimodal, ; nguyen2024multimodal, ; liu2025rna, ). The resulting representations are then fused through techniques such as attention mechanisms, contrastive learning, or embedding concatenation. This multi-branch strategy allows models to learn complementary features from both modalities while preserving the unique contributions of each, leading to enhanced representation quality and generalizability.
5.2. Integration of Sequence and Functional Representations
@@
While Protein Language Models (PLMs) have advanced the understanding of sequence patterns, they often lack the explicit biological knowledge required to accurately capture protein function. Sequence data alone provide limited information on a protein’s specific roles or its molecular interactions within cellular environments, which can hinder model performance in tasks such as prediction of protein functions, protein contact mapping, and protein-protein interaction analysis. Given that a protein’s structure fundamentally determines its function, models can be significantly improved by incorporating functional knowledge that highlights similarities among proteins with comparable shapes or sequences. Functional annotations provide comprehensive insights into the biological processes, molecular functions, and cellular components of a protein. Integrating such functional data into PLMs enriches protein representations, allowing models to more effectively generalize across diverse biological tasks and applications.
@@
To address the limitations of sequence-only models, functional data are typically treated as a separate modality. Researchers employ biomedical language models such as BioBERT (lee2020biobert, ) and PubMedBERT (gu2021domain, ), which are trained in large-scale biomedical literature and ontologies, to generate functional embeddings that capture rich contextual and domain-specific knowledge. These functional embeddings are then fused with sequence-derived representations using methods such as contrastive learning or joint embedding frameworks (zhang2022ontoprotein, ; zhou2023protein, ; xu2023protst, ; kilgore2025protein, ). This multimodal integration enhances the representational capacity of PLMs, enabling more accurate and biologically informed predictions on a range of functional and interaction-based tasks.
5.3. Unified Representations of Sequence, Structure, and Function
@@
Unified representations are used to simultaneously integrate sequence, structural, and functional data into a cohesive modeling framework, providing a comprehensive view of protein characteristics. Although sequence-based models capture evolutionary and contextual relationships and structure-based models provide insights into three-dimensional conformation, the integration of functional annotations adds a critical biological context regarding the roles and interactions of a protein within cellular systems. By combining these modalities, unified representations offer a holistic understanding of proteins, allowing models to capture the complex interaction between a protein’s linear sequence, its folded structure, and its biological function.
@@
To achieve this, researchers typically adopt multi-branch architectures, where each modality (sequence, structure, and function) is encoded separately using specialized models. The output of these models is then fused using techniques such as attention mechanisms, contrastive learning, or cross-modal embedding fusion (hu2023multimodal, ; abdine2024prot2text, ). This approach ensures that the unique contributions of each modality are preserved while enabling the model to learn complementary features that improve the overall quality of the representation. By capturing the full complexity of protein biology, these multimodal approaches pave the way for more accurate, generalizable, and biologically meaningful models.
6. Complex-Based Approaches
@@
6.1. Protein-Ligand Complex Representation
@@
The protein-ligand complex representation aims to learn a unified representation that effectively captures the interactions between proteins and small molecules, enabling the models to better understand binding mechanisms and affinity relationships. Rather than representing the entire protein structure, these approaches typically focus on the binding pocket, the localized region where molecular interactions occur. Upon concentration on this functionally relevant site, models can capture more accurately the spatial, chemical, and sequence-dependent features that govern ligand binding and molecular recognition. Approaches for protein-ligand complex representation can be broadly classified into two following categories, differentiated by their use of 3D structural information.
@@
The first category follows a dual-encoder framework (karimi2019deepaffinity, ; wang2022electra, ; pei2022smt, ; gao2023co, ; nakata2023end, ; nguyen2024proteinrediff, ; gao2024drugclip, ), where proteins and ligands are encoded separately, without explicitly modeling their spatial interactions. Protein representations are commonly derived from protein language models (PLMs) or graph-based encoders, while ligand representations are generated using SMILES-based transformers, molecular fingerprints, or graph neural networks (GNNs). Since these models do not require pre-existing 3D structural data for the full complex, they are particularly advantageous for high-throughput virtual screening and cases where structural information is unavailable. During fine-tuning, a lightweight interaction module is introduced to align and refine these representations for downstream tasks such as binding affinity prediction and molecular docking. This modular approach improves scalability and facilitates transfer learning across diverse protein-ligand datasets.
@@
The second category focuses on fine-grained interaction modeling, where 3D structural information of the complete protein-ligand complex is explicitly incorporated to learn atomic- or residue-level interactions. Unlike dual-encoder approaches, which align representations post hoc, these methods directly encode spatial and chemical dependencies within the binding pocket, capturing fine-grained molecular interactions. The binding mechanism of protein-ligand complexes is highly intricate, involving a diverse array of non-covalent forces such as $\pi$-stacking, $\pi$-cation interactions, salt bridges, water bridges, hydrogen bonds, hydrophobic interactions, and halogen bonds (C7MD00381A, ). Previous studies (zheng2019onionnet, ; moon2022pignet, ; feng2024proteinligand, ) have demonstrated that explicitly modeling these atomic-level forces significantly improves predictive performance, underscoring the importance of 3D molecular force modeling in PRL.
6.2. Protein Complex Representation
@@
Protein complexes consisting of two or more interacting polypeptide chains are fundamental to biological processes such as signal transduction, immune response, enzymatic activity, and molecular assembly. These complexes exhibit significant diversity in size, stability, and interaction mechanisms, ranging from small transient interactions to large, stable macromolecular structures such as ribosomes, chaperones, and antibody-antigen complexes. The representation of protein complexes closely follows the methodology used in protein-ligand complex modeling, where the goal is to learn a joint representation that captures both individual protein properties and interaction-specific features. Approaches to this problem can be broadly categorized on the basis of whether they explicitly encode the full complex structure or rely on separate representations of individual proteins.
@@
Structure-free approaches, which encode individual proteins separately before aligning them in a shared representation space, are widely applied to various challenges in protein complex modeling (yuan2023dg, ; li2024mvsf, ; bandara2024deep, ). These approaches do not explicitly model the structural constraints of the full complex during encoding. Instead, interactions are inferred post hoc through mechanisms such as contrastive learning or learned scoring functions. Among the many tasks tackled by these approaches, the prediction of the structure of the protein complex, the determination of the spatial arrangement of the interacting polypeptides given only sequence or monomeric structural information, is one of the most critical (evans2021protein, ; giulini2023towards, ; ruffolo2023fast, ).
@@
Structure-aware approaches, on the other hand, explicitly encode the entire 3D structure of the protein complex, incorporating spatial and physicochemical constraints to model binding interfaces and interaction dependencies (reau2023deeprank, ; yue2024integration, ; wang2024eggnet, ). These methods leverage geometric deep learning, graph-based modeling, and energy-based optimization techniques to refine complex formation by capturing atomic-level interactions and physical constraints. Because these approaches explicitly encode the joint structural context of interacting proteins, they can provide more interpretable predictions, improve docking accuracy, and enhance the design of protein-protein interactions for therapeutic applications. However, they require high-quality structural data, which may not always be available, and must balance computational efficiency with accuracy when handling large macromolecular assemblies.
Table 1. Comparative Summary of Strengths and Limitations of PRL Approaches.
@@
7. Databases for Protein Representation Learning
Large-scale biological databases are essential for Protein Representation Learning (PRL), which offers the sequence, structure, and functional data required for model development and evaluation. These resources have not only enabled large-scale pretraining, but have also shaped the design and evaluation of diverse representation methods. Table 2 summarizes key databases in different data modalities, reflecting their scale and utility in modern PRL research. Despite the breadth of available data, there are significant imbalances between modalities. Sequence data are highly abundant, with repositories such as UniProtKB and MGnify containing hundreds of millions to billions of entries, providing a rich foundation for capturing evolutionary and biochemical patterns in proteins. In contrast, experimentally resolved protein structures remain relatively limited, as databases such as the Protein Data Bank (PDB) are constrained by the time, cost, and complexity of structural determination techniques. The emergence of structure prediction resources, including AlphaFold DB and ESMAtlas, has significantly increased structural coverage, supporting the development of structure-aware models. However, these predicted models may vary in accuracy and often lack experimental validation, particularly in regions of intrinsic disorder or low confidence. Functional annotations represent an even sparser modality: Only a small subset of known proteins are associated with experimentally confirmed functions, and many annotations are computationally inferred, introducing varying degrees of uncertainty. These imbalances pose a particular challenge for multimodal approaches that rely on aligned and balanced data across sequence, structure, and function.